OpenStudy (anonymous):

Determine the slop off the tangent at x=4 for f(x)=x/(x^2 -15). would be help full if is someone can explain it using Limits.

OpenStudy (jdoe0001):

looks to me like it's asking for what is the limit of f(x) as x -> 4, approaches 4

OpenStudy (jdoe0001):

the value will be the slope as x = 4, and thus the limit

OpenStudy (anonymous):

no i have to use f(a+h)= f(a+h)-f(a) all over h

OpenStudy (anonymous):

so ur given 4 and when u plug that in to the function u get a value and that one of ur points on the graph. and u need to find another point basically so use the limit equation as i mention above

OpenStudy (jdoe0001):

right, still the same thing

OpenStudy (jdoe0001):

f(a+h)= f(a+h)-f(a) all over h <--- all it does is finds the derivative, which is the equation of the slope of the tangent line

OpenStudy (anonymous):

but i am having difficulties getting rid of the h in the denominator, and brb

OpenStudy (anonymous):

k back

OpenStudy (jdoe0001):

ever got rid of it?